OpenStudy (anonymous):

If 2tanx/(1-tan^2x)=1/sqrt3 then x can equal: A) pi/12 B) 7pi/12 C) 11pi/12 D) 5pi/12 Each answer includes nPi multiples, and I understand that tan=pi/3 or tan60, but the values in the answer choice are confusing me.

OpenStudy (anonymous):

tan(2x)=tan(pi/6) x=pi/12

OpenStudy (anonymous):

Why is that? I forgot to say it's multiple answer.

OpenStudy (mertsj):

\[\frac{2\tan x}{1-\tan ^2x}=\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

the answer is A

OpenStudy (mertsj):

\[2\sqrt{3}\tan x=1-\tan ^2x\]

OpenStudy (mertsj):

\[\tan ^2x+2\sqrt{3}\tan x-1=0\]

OpenStudy (debbieg):

2tanx/(1-tan^2x)=tan(2x) That's the double-angle Identity for tangent. So the equation is: \[\tan(2x)=\frac{ 1 }{\sqrt{3} }\] So now you need to solve with methods for multiple angle equations... do you know what I mean by that?

OpenStudy (anonymous):

@DebbieG That's where I get stuck. I only know how to get the base answers, not the multiple angles.

OpenStudy (debbieg):

Also, part of the problem might be that you say above: "tan=pi/3" By which I presume you mean to say that \[\tan(\frac{ \pi }{ 3 })=\frac{ 1 }{\sqrt{3} }\]....... you might want to double check that?? :)

OpenStudy (debbieg):

OK, here's the key with MA equations: PUT BLINDERS ON. :) In other words, pay no attention to the fact that it's a multiple angle just yet. We'll get to that soon enough. Just supposed you had \[\tan(x)=\frac{ 1 }{\sqrt{3} }\] What would be the solution? (in radians, since your answers are in radians)

OpenStudy (anonymous):

I would think it was \[\pi/3\] but my foundation of this isn't too swell.

OpenStudy (debbieg):

You should know this triangle like the back of your hand. :) And although I did it in degrees, obviously it works for the equivalent radian measures. 

OpenStudy (anonymous):

So wait... Is it pi/6? I said pi/3 because I divided sin by cos for that point on the unit circle and thought it was 1/sqrt3... I'm going about this the wrong way, huh? :)

OpenStudy (debbieg):

Now, that would work, but sin(pi/6)=1/2, cos(pi/6)=sqrt(3)/2 They are opposite for pi/3, so you get tan(pi/3)=sqrt(3) Personally I prefer the triangle, but the unit circle will work if you use the right values. :)

OpenStudy (debbieg):

OK, so if we have an equation like\[\tan(x)=\frac{ 1 }{\sqrt{3} }\]then we know that \[x=\frac{ \pi }{6 }\]is one solution. But here, it isn't x that is in he function, it's 2x... right? So that means the solution we found is a solution for 2x, not for x. That is: \[2x=\frac{ \pi }{ 6 }\]right? Understand so far?

OpenStudy (debbieg):

Now as you said, we need to add on the multiples of the period, to capture all the coterminal solutions. Now is the time for that: \[\large 2x=\frac{ \pi }{ 6 }+k\pi\]

OpenStudy (debbieg):

Now all that's left is to solve THIS for x.... but don't forget to distribute on the RHS when you do so! Thus, the fact that it's a multiple angle will affect not just the solution, but the increment between solutions - the kpi. Follow me? So you said above that "Each answer includes nPi" and that's normally true for a tangent equation... but since it's a multiple angle, it will NOT be nPi here when you're all finished.

OpenStudy (anonymous):

So multiply pi/6 times 2? I'm having difficulty getting the numbers in front pi in the answers. (Ex: The '5' in 5pi/12). How do any relate to pi/6?

OpenStudy (debbieg):

Look at this again: \[\large 2x=\frac{ \pi }{ 6 }+k\pi\] What do you need to do, in order to solve it for x? Tell me what the operation is....

OpenStudy (anonymous):

Oops, wait... Divide by 2?

OpenStudy (debbieg):

There ya go... :)

OpenStudy (anonymous):

Phew, online class overload, sorry. :) So, okay. Pi/12. Are there others? How are those calculated?

OpenStudy (debbieg):

Just the ones that come from the periodicity of the function. That's what I was talking about above - but those are captured by the +kPi, which is now +kPi/2, because of the division due to the multiple angle. Follow? That's why I said early on, "put blinders on" - you need to forget about the multiple angle, until you have the whole "general form" solution, including the +kPi (I call it "the incrementor" since it accounts for the increment between solutions, but that's my own word, don't like for it in a trig book, lol). So if this wasn't a multiple choice question, you would get as your complete solution set: \[\left\{ x|x=\frac{ \pi }{ 12 }+k \cdot\frac{ \pi }{ 2 } \right\}\]or a bit less formally, just:\[x=\frac{ \pi }{ 12 }+k \cdot\frac{ \pi }{ 2 } \]

OpenStudy (anonymous):

Okay, I think I finally understand up to that point.

OpenStudy (debbieg):

I should add, IF this was a sine or cosine multiple angle equation, you WOULD have an additional piece to the solution set, because there would be a second solution that you have to find. Then add the incrementor to both. The difference here is that tangent has a period of pi, rather than 2pi for sine and cosine, so you only need the 1 solution, the incrementor takes care of the rest of it.

OpenStudy (anonymous):

I remember learning sine and cosine, but this course never went over the more complex side of tan. That's why I'm kinda clueless for this equation.